OpenStudy (anonymous):

A vector v forms with axis 0x and 0y the angles alpha=120º and beta=45º. What angle does v make with 0z? i know the answer is gama=45º, but how do i get there?

OpenStudy (anonymous):

is anyone gonna answer this? i really need it.

OpenStudy (dumbcow):

having little trouble working it out...hold on

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

just leave it, thanks for the time.. :)

OpenStudy (anonymous):

fooound it. cos^2alpha+cos^2beta+co^2gama=1

OpenStudy (anonymous):

\[\cos^2(\gamma) =1 - \cos^2(\alpha) - \cos^2(\beta) =\\ 1- \frac 1 4 - \frac 1 2=\frac 14\\ \beta = 30 \]